Calculate the number of moles and molarity of n/2 500 ml solution of H2SO4


1
Expert's answer
2021-06-25T05:41:00-0400

See,

Normality=Molarity*n-factor

For H2SO4 n factor is 2 due to presence of two replacable H+ ions

Given Normality is 0.5N

Hence Molarity will be 0.5/2=0.25M

Number of moles=Molarity*Volume(in L)

Moles= 0.25*500/1000= 0.125